OpenStudy (anonymous):

Choose the pair of events that is mutually exclusive. (a). Even integers and multiples of 3 (b). Odd integers and multiples of 4 (c). Integers greater then 40 and integers less then 50. (d). Prime numbers and even integers

ganeshie8 (ganeshie8):

Mutually exclusive events can't happen at the same time

ganeshie8 (ganeshie8):

look at option a : are there any even numbers in multiples of 3 ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

think again : 6 = 3*2, which is an even number and also a multiple of 3, right ?

OpenStudy (anonymous):

right....

OpenStudy (anonymous):

would 12 be one too?

ganeshie8 (ganeshie8):

Exactly ! so can we safely eliminate option A ? cuz "even numbers" and "multiples of 3" are not mutually exclusive.

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

what about option B ?

OpenStudy (anonymous):

ummm Maybe b is the answer becasue i cant think of anything hang on

OpenStudy (anonymous):

theres no odd numbers in multipples of four

ganeshie8 (ganeshie8):

Excellent !!

ganeshie8 (ganeshie8):

"odd integers" and "multiples of 4" are mutually exclusive !

OpenStudy (anonymous):

:) Thank you

ganeshie8 (ganeshie8):

yw :)

OpenStudy (anonymous):

if P(Q)=3/4 and P(R)=1/6 find P(Q or R)